# PPNet: A Novel Neural Network Structure for End-to-End Near-Optimal Path Planning

### v1

@@ Abstract
The classical path planners, such as sampling-based path planners, have the limitations of sensitivity to the initial solution and slow convergence to the optimal solution. But finding a near-optimal solution in a short period of time is challenging in many applications such as the autonomous vehicle with limited power/fuel. For achieving an end-to-end near-optimal path planner, we first divide the path planning problem into two subproblems, which are path’s space segmentation and waypoints generation in the given path’s space. And we further propose a two-level cascade neural network named Path Planning Network (PPNet) to solve the path planning problem by solving the abovementioned subproblems. Moreover, we propose a novel efficient data generation method for path planning named EDaGe-PP. The results show the total computation time is less than 1/33 and the success rate of PPNet trained by the dataset that is generated by EDaGe-PP is about 2×2$\times$2 × compared to other methods. We validate PPNet against state-of-the-art path planning methods. The results show PPNet can find a near-optimal solution in 15.3ms, which is much shorter than the state-of-the-art path planners.
@@
Deep Learning, Near-Optimal, End-to-End, Path Planning.
@@ I Introduction
Path planning is one of the core research problems in robotics. It is to find a collision-free, low-cost path connecting the initial state and goal state. The popular methods, such as RRT* [1] and its variants, can efficiently explore the state space. But due to the sensitivity to the initial solution and slow convergence to the optimal solution, they can’t guarantee to find a near-optimal solution within a specific time, especially when the required planning time is quite short in many applications such as the autonomous vehicle with limited power/fuel. Thus many variants of RRT* were proposed to shorten computation time. For example, Informed RRT* (IRRT*) [2] is a representative variant of implementing Informed Search on RRT*. After finding the initial solution, IRRT* can converge to the optimal solution more quickly. Batch Informed Trees (BIT*) [3], and Advance BIT* (ABIT*) [4] are to shorten computation time by further using the heuristics function to direct the process of exploring state space. It makes the methods find the initial solution more quickly in some cases. However, the heuristic sampling domain of the abovementioned methods might be quite large in some cases, and sampling in the domain remains random in these methods. The inefficiency of exploring the state space makes the computation time of the methods might be unacceptable for the applications mentioned above.
@@
Recently, machine learning techniques are used in path planning for overcoming the limitations of sampling-based path planners. Many learning-based methods can perform quite short computation time in the learned environments. Neural RRT* (NRRT*) [5] combines the machine learning techniques and the sampling-based method. NRRT* uses a neural network for segmenting the promising region in the state space. Letting RRT* sample in the segmented region, which is smaller than the sampling domain of the methods used informed search, can accelerate the process of finding the initial path and converging to the optimal path. Motion Planning Networks (MPNet) [6] uses RNN for predicting the next waypoint directly. Without random sampling, MPNet can lay out a near-optimal path with the necessary iterations in some cases. But these works still need increase iterations for improving the success rate, which makes computation time uncontrollable. To overcome this limitation, we think that developing an end-to-end near-optimal path planner is a promising way, which can provide theoretical guarantees for finding a near-optimal path in a short period of time.
@@
To achieve a method that can find a near-optimal solution in an end-to-end way, we first divide the path planning problem into two subproblems, which are path’s space segmentation and waypoints generation in the given path’s space. And we further propose a novel two-level cascade neural network named Path Planning Network (PPNet) that can solve the path planning problem by solving the two subproblems orderly. PPNet comprises a segmentation module as the first level named SegNet and a generation module as the second level named GenNet. SegNet takes the map with the initial point and goal point mark as input and outputs the path space. GenNet takes the output of SegNet as input and outputs the probability map indicating the probability of whether each point in the environment is the waypoint. With the probability map, the path can be extracted based on a simple rule during the computation time of PPNet. Moreover, to achieve a better success rate of PPNet, we also propose a novel efficient data generation method for path planning named EDaGe-PP. The results show that EDaGe-PP provides about $\times3$ computation speed improvement and 2×2$\times$2 × success rate improvement compared with the popular methods. The data generated by EDaGe-PP and the two-level cascade structure of PPNet enable PPNet to achieve about $41\%$ success rate improvement compared with the representative learning-based method and find a near-optimal path in 15.3ms (Fig. 1).
@@ II Related Work
The performance requirement that some applications such as the autonomous vehicle require a near-optimal path in a specific time is challenging to the computational efficiency of path planners. But due to the randomly ordered search, popular methods such as sampling-based methods (e.g., RRT* and PRM*) might not satisfy such a performance requirement. Two reasons for the inefficiency of randomly ordered search are random sampling in the whole state space, and expanding of the tree that is guided by the random samples. To improve the efficiency of expanding the tree, Fast Marching Tree (FMT*) [7] uses a marching method to process samples. But for achieving a 100% success rate of FMT*, the number of samples should be larger than 1000 and the execution time might be longer than 1s [7]. If the number of samples is less than that, the success rate will decrease with the number of samples rapidly. The methods used Informed search [2] sample in the heuristic sampling domain, which is determined by a hypothesis or information that has been known. Instead of sampling in the whole state space, the methods improve the convergence to the optimal path. But the sampling domain might be quite large in some cases, and the expanding of the tree is still inefficient. BIT* [3] further uses incremental graph-search techniques like Lifelong Planning A* (LPA*) [8] to guide the expanding of the tree. ABIT* [4] achieves a shorter time of finding an initial path by using more advanced graph-search techniques, such as inflation and truncation. It means BIT* and ABIT* are to shorten the computation time by improving both the sampling and the expanding. But these improvements might be insufficient for some applications because the heuristic function that directs the expanding of the tree makes these methods might get trapped in some cases.
@@
By applying the machine learning techniques there are methods that achieve ordered search. NEED [9] uses a segmentation model to segment the promising region of the state space as the preprocessing for sampling-based methods. The promising region is the region that has a high probability that waypoints are in it. Neural RRT* [5] and 3D Neural RRT* [10] can be seen as the application of NEED which combine RRT* with NEED. Neural RRT* and 3D Neural RRT* improve the efficiency of search by searching in the promising region which can be seen as the variant of informed search. But instead of searching in the ellipse that is drawn by the cost of the initial solution, they search in the promising region which is segmented by the segmentation model. The more narrow sampling region makes the sampling more efficient. MPNet [6] avoids random sampling and expanding by using a recurrent neural network (RNN) to output the next waypoint directly. And due to bidirectional planning like RRT-Connect [11], MPNet alleviates the difficulty of predicting the long sequence of RNN. MPNet could lay out the near-optimal solution in some learned environments without generating extra nodes. But in complex environments the necessary number of iterations of the abovementioned methods might be quite large, and for improving the quality of the path the number needs to be further increased.
@@
Unlike the previous works that apply machine learning techniques, PPNet solves the path planning problem by solving the two subproblems orderly, which makes PPNet can output a path in an end-to-end way. Moreover, the continuous-curvature paths with analytical expression, which is generated by EDaGe-PP, make PPNet achieve more efficient training and a much higher success rate.
@@ III EDaGe-PP: Efficient Data Generation
In this section, we describe EDaGe-PP. The popular data generation methods used in the works that try to apply deep learning techniques to path planning are to use classical planners such as RRT* for finding an optimal path. But these data generation methods have several limitations. Firstly, using classical planners to generate an optimal path is extremely time-consuming in complex environments. Secondly, the paths only contain a few waypoints is computationally expensive to generate data with the custom formats for the training of neural networks, which is usual in developing learning-based methods. Thirdly, they both need appropriate a prior discretization for the path planning problem with clearance requirements, which is difficult sometimes. Fourthly, the paths comprised of discrete waypoints make it difficult for the neural networks to converge. For overcoming the limitations of popular data generation methods mentioned above, we propose an efficient data generation method for path planning named EDaGe-PP. EDaGe-PP can generate data with analytic paths while satisfying the clearance requirement. Moreover, EDaGe-PP provides about 33× computation speed improvement and 2× success rate improvement compared with the popular methods.
@@ III-A Random Path Generation
In this paper, we use a predetermined number of polynomial functions to describe the paths which are randomly generated by our algorithm. The polynomial function is defined as follows:
$y=w^{\mathrm{T}}\cdot\vec{X},\$ (1)
$\vec{w}=\left[w_{0},w_{1},\cdots,w_{r}\right]^{\mathrm{T}},\$ (2)
$\vec{X}=\left[1,x,x^{2},\cdots,x^{r}\right]^{\mathrm{T}},\$ (3)
@@ III-B Space of the Path
where $n^{{}^{\prime}}$ denotes the number of the samples of the boundary around the initial point or the end point.
@@ III-C Settings of the Obstacles
For making the path a near-optimal solution in any state space particular constraints need to be applied. In this paper, we use obstacles that are circular in shape as the constraints for the path. The calculation of settings of the obstacles is presented in Alg. 3. The process of calculating the settings of obstacles is divided into two steps. The first step is to calculate the convex hull of the space and extract the subspace of the space. The second step is to calculate the size and position of obstacles in the subspace.
@@ III-D Random State Space
With the mask of the path’s space and corresponding settings of obstacles, we can generate the path planning problems that the target path is a near-optimal solution in them. The Algorithm for generating Random State Space is presented in Alg. 4.
The translation and rotation of the path are generated randomly. Then the same translation and rotation are applied to the convex hull of the path’s waypoints. Whether the convex hull of the path exceeds the boundary of the state space $\mathscr{R}$ is checked. If the convex hull is in the state space, the algorithm will continue executing, or (Alg. 4, Lines 44) will be continually repeated until the success of boundary check or the times of repetition exceeds the predetermined threshold. When the times of repetition have exceeded the threshold, the algorithm of path generation will be judged as a failure and the algorithm will be executed from the start (Alg. 1, Line 1). After boundary check, the size and the position of the circle obstacles are randomly generated and only the collision-free obstacles will be remained (Alg. 4, Lines 44). The settings of obstacles of the path are placed with the same translation and rotation as the path. The original mask representing the waypoints of the path, and the original mask representing the path’s space are doing the same translation and rotation with the path (Alg. 4, Lines 44). The image of the state space $I_{Prob}$ and the mask of the path’s space $I_{S}$ do Hadamard product for making sure that no obstacles in the path’s space, which means the path is collision-free. The calculation is described as:
@@ IV PPNet: End-to-End Path Planner
In this section, we describe the two-level cascade structure of PPNet. For enabling PPNet to solve path planning problems in an end-to-end way, we divide the path planning problem into two subproblems. The first one is to segment the path’s space where the path is in it while satisfying the clearance requirement. The second one is to generate the waypoints of the path based on the shape of the path’s space segmented in the first subproblem. This leads the structure of PPNet to a two-level cascade structure that each level solves one of the two subproblems (Fig. 3).
In Section IV-A, we introduce the first subproblem which is space segmentation, and the design of the corresponding first-level model of PPNet named SegNet. In Section IV-B, the second subproblem which is waypoints generation, and the design of the second-level model of PPNet named GenNet are introduced.
Figure 3: Model overview. PPNet consist of SegNet and GenNet. SegNet can solve the first subproblem, which is space segmentation. GenNet can solve the second subproblem, which is waypoints generation.
@@ IV-A Space Segmentation
Space segmentation is the task to segment the space of the path for the corresponding path planning problem. The segmented path’s space should be the space within the boundary of the path that is calculated in Section III-B. This makes the shape of the space can be used for generating the corresponding path. Such a problem can be described as:
@@
For making it easier to solve this subproblem by neural networks, we further transfer the problem into a binary classification task, which can be defined as:
@@
where $\mathcal{S}$ denotes the function of SegNet, $\mathcal{M}$ denotes the picture that describes the path planning problems, and $Y$ denotes the result of space segmentation.
This leads the problem to a segmentation task, so we develop a segmentation model to solve the first subproblem which is space segmentation, named SegNet. SegNet takes the picture that describes the path planning problem as the input and outputs the mask of the path’s space. The output of SegNet is a mask of the path’s space which is a binary picture that indicates whether the point is in the path’s space, which means each pixel is the result of the problem defined in Equation 37 of the corresponding point $x$. And we use such a picture as the representation of the path’s space.
@@ IV-B Waypoints Generation
Waypoints generation is the task to generate the waypoints by the shape of the path’s space. The problem can be described as:
$W=f_{G}\left(S\right),\$ (38)
@@
$M=\mathcal{G}\left(Y\right),\$ (39)
@@
The first step can be seen as a generation task, so we develop a generation model to solve it, named GenNet. GenNet is designed as a Transformer-based Autoencoder. GenNet takes the binary picture $Y$ that represents the path’s space as input and outputs the probability map of the waypoints.
The second step is to extract the waypoints from the probability map $M$. $\mathcal{E}\left(\cdot\right)$ denotes the simple rule of extracting waypoints from the probability map that is generated by GenNet, which is defined as:
@@
where $X_{neighbor}$ is the neighborhood points of the current point on the probability map, which starts at the initial point. The computation of extracting the waypoints from the probability map can be finished during the inference time of PPNet.

### v2

# PPNet: A Two-Stage Neural Network for End-to-end Path Planning

@@ Abstract
The classical path planners, such as sampling-based path planners, can provide probabilistic completeness guarantees in the sense that the probability that the planner fails to return a solution if one exists, decays to zero as the number of samples approaches infinity. However, finding a near-optimal feasible solution in a given period is challenging in many applications such as the autonomous vehicle. To achieve an end-to-end near-optimal path planner, we first divide the path planning problem into two subproblems, which are path space segmentation and waypoints generation in the given path’s space. We further propose a two-stage neural network named Path Planning Network (PPNet) each stage solves one of the subproblems abovementioned. Moreover, we propose a novel efficient data generation method for path planning named EDaGe-PP. EDaGe-PP can generate data with continuous-curvature paths with analytical expression while satisfying the clearance requirement. The results show the total computation time of generating random 2D path planning data is less than 1/33 and the success rate of PPNet trained by the dataset that is generated by EDaGe-PP is about 2×2$\times$2 × compared to other methods. We validate PPNet against state-of-the-art path planning methods. The results show that PPNet can find a near-optimal solution in 15.3ms, which is much shorter than the state-of-the-art path planners.
@@
Neural Networks, End-to-End, Path Planning.
@@ I Introduction
Path planning is one of the core research problems in robotics. It is to find a collision-free, low-cost path connecting the initial state and goal state. The popular methods, such as sampling-based path planners, can provide probabilistic completeness guarantees, which means the probability that the planners succeed in returning a solution increases by increasing the number of samples [1]. Therefore, they can’t guarantee to find a near-optimal solution within a specific time, especially when the required planning time is quite short (i.e., $\leq$0.1s) in many applications such as the autonomous vehicle. Thus many variants of RRT* were proposed to shorten computation time. For example, Informed RRT* (IRRT*) [2] is a representative variant of implementing Informed Search on RRT*. After finding the initial solution, IRRT* can converge to the optimal solution more quickly. Batch Informed Trees (BIT*) [3], and Advance BIT* (ABIT*) [4] are to shorten computation time by further using the heuristics function to direct the process of exploring state space. It makes the methods find the initial solution more quickly in some cases. However, the heuristic sampling domain of the abovementioned methods might be quite large in some cases, and sampling in the domain remains random in these methods. The inefficiency of exploring the state space makes the computation time of the methods might be unacceptable for the applications mentioned above.
@@
Recently, machine learning techniques have been used in path planning to overcome the limitations of sampling-based path planners. Many learning-based methods can perform quite short computation time in the learned environments. Neural RRT* (NRRT*) [5] combines the machine learning techniques and the sampling-based method. NRRT* uses a neural network for segmenting the promising region in the state space. Letting RRT* sample in the segmented region, which is smaller than the sampling domain of the methods used informed search, can accelerate the process of finding the initial path and converging to the optimal path. Motion Planning Networks (MPNet) [6] uses RNN for predicting the next waypoint directly. Without random sampling, MPNet can lay out a near-optimal path with the necessary iterations in some cases. However, these works still need to increase iterations to improve the success rate, which makes computation time uncontrollable.
@@
To overcome the limitation abovementioned, we think that developing an end-to-end near-optimal path planner is a promising way that provide theoretical guarantees for finding a near-optimal path in a short period. To achieve a method that can find a near-optimal solution in an end-to-end way, we first divide the path planning problem into two subproblems, which are path space segmentation and waypoints generation in the given path’s space. We further propose a novel two-stage neural network named Path Planning Network (PPNet) that can solve the path planning problem by solving the two subproblems orderly. PPNet comprises a segmentation model as the first stage named SpaceSegNet and a generation model as the second stage named WaypointGenNet. SpaceSegNet takes the map with the initial point and goal point mark as input and outputs the path space. WaypointGenNet takes the output of SpaceSegNet as input and outputs the probability map indicating the probability of whether each point in the environment is the waypoint. With the probability map, the path can be extracted based on a simple rule during the computation time of PPNet. Moreover, to achieve a better success rate of PPNet, we also propose a novel efficient data generation method for path planning named EDaGe-PP. The results show that EDaGe-PP provides about $\times3$ computation speed improvement and 2×2$\times$2 × success rate improvement compared with the popular methods. The data generated by EDaGe-PP and the two-stage structure of PPNet enable PPNet to achieve about $41\%$ success rate improvement compared with the representative learning-based method and find a near-optimal path in 15.3ms (Fig. 1).
@@ II Related Work
The performance requirement that some applications such as the autonomous vehicle require a near-optimal path in a specific time is challenging to the computational efficiency of path planners. However, due to the randomly ordered search, popular methods such as sampling-based methods (e.g., RRT* and PRM*) might not satisfy such a performance requirement. Two reasons for the inefficiency of randomly ordered search are random sampling in the whole state space and expansion of the tree that is guided by the random samples. To improve the efficiency of expanding the tree, Fast Marching Tree (FMT*) [7] uses a marching method to process samples. But for achieving a 100% success rate of FMT*, the number of samples should be larger than 1000 and the execution time might be longer than 1s [7]. If the number of samples is less than that, the success rate will decrease with the number of samples rapidly. The methods used Informed search [2] sample in the heuristic sampling domain, which is determined by a hypothesis or information that has been known. Instead of sampling in the whole state space, the methods improve the convergence to the optimal path. But the sampling domain might be quite large in some cases, and the expansion of the tree is still inefficient. BIT* [3] further uses incremental graph-search techniques like Lifelong Planning A* (LPA*) [8] to guide the expansion of the tree. ABIT* [4] achieves a shorter time of finding an initial path by using more advanced graph-search techniques, such as inflation and truncation. It means BIT* and ABIT* are to shorten the computation time by improving both the sampling and the expansion. However, these improvements might be insufficient for some applications because the heuristic function that directs the expanding of the tree makes these methods might get trapped in some cases.
@@
By applying the machine learning techniques there are methods that achieve ordered search. NEED [9] uses a segmentation model to segment the promising region of the state space as the preprocessing for sampling-based methods. The promising region is the region that has a high probability that waypoints are in it. Neural RRT* [5] and 3D Neural RRT* [10] can be seen as the application of NEED which combines RRT* with NEED. Neural RRT* and 3D Neural RRT* improve the efficiency of search by searching in the promising region which can be seen as the variant of informed search. But instead of searching in the ellipse that is drawn by the cost of the initial solution, they search in the promising region which is segmented by the segmentation model. The more narrow sampling region makes the sampling more efficient. MPNet [6] avoids random sampling and expanding by using a recurrent neural network (RNN) to output the next waypoint directly. And due to bidirectional planning like RRT-Connect [11], MPNet alleviates the difficulty of predicting the long sequence of RNN. MPNet could lay out the near-optimal solution in some learned environments without generating extra nodes. But in complex environments the necessary number of iterations of the abovementioned methods might be quite large, and for improving the quality of the path the number needs to be further increased.
@@
Unlike the previous works that apply machine learning techniques, PPNet solves the path planning problem by solving the two subproblems orderly, which makes PPNet able to output a path in an end-to-end way. Moreover, the continuous-curvature paths with analytical expression, which is generated by EDaGe-PP, make PPNet achieve more efficient training and a much higher success rate.
@@ III EDaGe-PP: Efficient Data Generation
In this section, we describe EDaGe-PP. EDaGe-PP was proposed for generating the custom data more efficiently for the training of learning-based path planning methods. It can overcome four limitations of the popular data generation methods for learning-based methods. Firstly, using classical planners to generate an optimal path is extremely time-consuming in complex environments. Secondly, the paths only contain a few waypoints is computationally expensive to generate data with the custom formats for the training of neural networks, which is usual in developing learning-based methods. Thirdly, they both need appropriate a prior discretization for the path planning problem with clearance requirements, which is difficult sometimes. Fourthly, the paths comprised of discrete waypoints make it difficult for the neural networks to converge. To overcome the limitations of popular data generation methods mentioned above, we propose an efficient data generation method for path planning named EDaGe-PP. EDaGe-PP can generate data with continuous-curvature paths with analytical expression while satisfying the clearance requirement. Moreover, EDaGe-PP provides about 33× computation speed improvement and 2× success rate improvement compared with the popular methods.
@@ III-A Random Path Generation
In this paper, we use a predetermined number of polynomial functions to describe the paths that are randomly generated by our algorithm. The polynomial function is defined as follows:
$y=w^{\mathrm{T}}\cdot\vec{X},$ (1)
$\vec{w}=\left[w_{0},w_{1},\cdots,w_{r}\right]^{\mathrm{T}},$ (2)
$\vec{X}=\left[1,x,x^{2},\cdots,x^{r}\right]^{\mathrm{T}},$ (3)
@@ III-B Space of the Path
where $n^{{}^{\prime}}$ denotes the number of samples of the boundary around the initial point or the end point.
@@ III-C Settings of the Obstacles
For making the path a near-optimal solution in any state space particular constraints need to be applied. In this paper, we use obstacles that are circular as the constraints for the path. The calculation of the obstacles’ settings is presented in Alg. 3. The process of calculating the settings of obstacles is divided into two steps. The first step is to calculate the convex hull of the space and extract the subspace of the space. The second step is to calculate the size and position of obstacles in the subspace.
@@ III-D Random State Space
With the mask of the path’s space and corresponding settings of obstacles, we can generate the path planning problems that the target path is a near-optimal solution. The Algorithm for generating Random State Space is presented in Alg. 4.
The translation and rotation of the path are generated randomly. Then the same translation and rotation are applied to the convex hull of the path’s waypoints. Whether the convex hull of the path exceeds the boundary of the state space $\mathscr{R}$ is checked. If the convex hull is in the state space, the algorithm will continue executing, or (Alg. 4, Lines 44) will be continually repeated until the success of boundary check or the times of repetition exceeds the predetermined threshold. When the times of repetition have exceeded the threshold, the algorithm of path generation will be judged as a failure and the algorithm will be executed from the start (Alg. 1, Line 1). After the boundary check, the size and the position of the circle obstacles are randomly generated and only the collision-free obstacles will remain (Alg. 4, Lines 44). The settings of obstacles of the path are placed with the same translation and rotation as the path. The original mask representing the waypoints of the path, and the original mask representing the path’s space are doing the same translation and rotation with the path (Alg. 4, Lines 44). The image of the state space $I_{Prob}$ and the mask of the path’s space $I_{S}$ do Hadamard’s product for making sure that no obstacles in the path’s space, which means the path is collision-free. The calculation is described as:
@@ IV PPNet: End-to-End Path Planner
In this section, we describe the two-stage structure of PPNet. To develop a method that can find a solution for end-to-end path planning while satisfying the clearance requirement, we found a key characteristic of solutions satisfying clearance requirements, which is there is a specific relationship between the path and its space, which means the path can be obtained by the shape of its space. Moreover, while the path can satisfy the clearance requirement, the space of the path should be completely in the free space. Therefore, we divide the path planning problem into two subproblems. The first one is to segment the path’s space where the path is in it while satisfying the clearance requirement. The second one is to generate the waypoints of the path based on the shape of the path’s space segmented in the first subproblem. This leads the structure of PPNet to a two-stage structure in which each stage solves one of the two subproblems (Fig. 3).
In Section IV-A, we introduce the first subproblem which is space segmentation, and the design of the corresponding first-stage model of PPNet named SpaceSegNet. In Section IV-B, the second subproblem which is waypoints generation, and the design of the second-stage model of PPNet named WaypointGenNet are introduced.
Figure 3: Model overview. PPNet consists of SpaceSegNet and WaypointGenNet. SpaceSegNet can solve the first subproblem, which is space segmentation. WaypointGenNet can solve the second subproblem, which is waypoints generation.
@@ IV-A Space Segmentation
Space segmentation is the task of segmenting the space of the path for the corresponding path planning problem. The segmented path’s space should be the space within the boundary of the path that is calculated in Section III-B. This makes the shape of the space can be used for generating the corresponding path. Such a problem can be described as:
@@
To make it easier to solve this subproblem by neural networks, we further transfer the problem into a binary classification task, which can be defined as:
@@
where $\mathcal{S}$ denotes the function of SpaceSegNet, $\mathcal{M}$ denotes the picture that describes the path planning problems, and $Y$ denotes the result of space segmentation.
This leads the problem to a segmentation task, so we develop a segmentation model to solve the first subproblem which is space segmentation, named SpaceSegNet. SpaceSegNet takes the picture that describes the path planning problem as the input and outputs the mask of the path’s space. The output of SpaceSegNet is a mask of the path’s space which is a binary picture that indicates whether the point is in the path’s space, which means each pixel is the result of the problem defined in Equation 37 of the corresponding point $x$. And we use such a picture as the representation of the path’s space.
@@ IV-B Waypoints Generation
Waypoints generation is the task of generating the waypoints by the shape of the path’s space. The problem can be described as:
$W=f_{G}\left(S\right),$ (38)
@@
$M=\mathcal{G}\left(Y\right),$ (39)
@@
The first step can be seen as a generation task, so we develop a generation model to solve it, named WaypointGenNet. WaypointGenNet is designed as a Transformer-based Autoencoder. WaypointGenNet takes the binary picture $Y$ that represents the path’s space as input and outputs the probability map of the waypoints.
The second step is to extract the waypoints from the probability map $M$. $\mathcal{E}\left(\cdot\right)$ denotes the simple rule of extracting waypoints from the probability map that is generated by WaypointGenNet, which is defined as:
@@
where $X_{neighbor}$ is the neighborhood points of the current point on the probability map, which starts at the initial point. The probability map $M$ can be seen as the solution to the given path planning problem. The computation of extracting the waypoints from the probability map can be finished during the executed time. So the total computation time of our method is defined as:
$t=t_{\mathcal{S}}+t_{\mathcal{G}},$ (42)
where $t_{\mathcal{S}}$ is the inference time of SpaceSegNet $\mathcal{S}$, $t_{\mathcal{G}}$ is the inference time of WaypointGenNet $\mathcal{G}$.